OpenStudy (anonymous):

the series ((-1)^n+1)(x-2)^n)/n .... the interval of convergence is 1

OpenStudy (anonymous):

another end point one, this time i do it with paper before typing

OpenStudy (anonymous):

\[|x-2|<1\]

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

now we try replacing x by 1 btw you have done all the hard work, the rest should be the easy part

OpenStudy (anonymous):

at x=1 i think it is \[\sum_{?}^{?} (-1)^{n+1}(-1)^{n}\div n\] ???

OpenStudy (saifoo.khan):

Whenever free, http://openstudy.com/study#/updates/4f333c8be4b0fc0c1a0b72dd

OpenStudy (anonymous):

ok very smilar to the last one

OpenStudy (anonymous):

but this time we don't do anything stupid \[(-1)^{n+1}(-1)^n=(-1)^{2n+1}=-1\]

OpenStudy (anonymous):

on account of \[2n+1\] is odd.

OpenStudy (anonymous):

so does not converge at x = 1

OpenStudy (anonymous):

what do you want to bet it does converge at x = 3?

OpenStudy (anonymous):

.ah! um a lot? haha.

OpenStudy (anonymous):

yeah well we can just about visualize it because it is the same as the last one. this time you get \[(-1)^{n+1}\times 1=(-1)^{n+1}\]so we are in good shape

OpenStudy (anonymous):

thank you! another question...

OpenStudy (anonymous):

this one alternates in other words, whereas the other one does not

OpenStudy (anonymous):

keep posting

OpenStudy (anonymous):

you got zarkon watching so if i mess up he will step in wish he had stepped in to the last one so i wouldn't have made a fool out of myself